# Look the Other Way: Designing ‘Positive’ Molecules with Negative Data via Task Arithmetic

### v1

@@ Abstract
The scarcity of molecules with desirable properties (i.e., ‘positive’ molecules) is an inherent bottleneck for generative molecule design. To sidestep such obstacle, here we propose molecular task arithmetic: training a model on diverse and abundant negative examples to learn ‘property directions’ – without accessing any positively labeled data – and moving models in the opposite property directions to generate positive molecules. When analyzed on 20 zero-shot design experiments, molecular task arithmetic generated more diverse and successful designs than models trained on positive molecules. Moreover, we employed molecular task arithmetic in dual-objective and few-shot design tasks. We find that molecular task arithmetic can consistently increase the diversity of designs while maintaining desirable design properties. With its simplicity, data efficiency, and performance, molecular task arithmetic bears the potential to become the de-facto transfer learning strategy for de novo molecule design.
@@ 1 Introduction
Discovering one drug molecule can take over a decade and billions of dollars1, 2. The first obstacle is charting the ‘chemical space’3 effectively. Chemical space is estimated to contain approximately 1060 drug-like molecules, with scarcity of molecules possessing desirable properties, e.g., bioactivity towards a pharmaceutically relevant target. Generative deep learning has emerged as a revolutionizing technology for drug discovery in the last decade – with the potential to shorten de novo design pipelines from years to weeks 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16.
Transfer learning has been fundamental for generative deep learning in molecule design17. The generative models are first pretrained on large, unlabeled molecule datasets to learn unconditional molecule generation. Next, the pretrained models are fine-tuned using molecules of a target property (high solubility, bioactivity, synthetic accessibility, etc.) and sampled to design compounds satisfying the design task18, 19.
The fine-tuning stage faces an inherent challenge: molecules with desirable properties are not only scarce but also lack structural diversity20. In the case of early stage drug discovery, positive molecules (or ‘hits’) can be as rare as 1% in large, diverse molecule libraries. These molecules are often optimized with small structural edits to explore structure-activity relationships, resulting datasets typically containing a few hundred molecules with low structural diversity21, 22, 23, 24. Here, we target such key data limitations in molecule design by proposing a new strategy for generative deep learning – molecular task arithmetic.
Molecular task arithmetic targets the limited availability of desirable and structurally diverse molecules in drug discovery. By identifying directions in the model weight space via training on abundant and diverse negative molecules, molecular task arithmetic moves pretrained model weights in the desirable property direction to create positive molecules. Editing models with task directions is known as task arithmetic25, and has already powered successful applications in computer vision and natural language processing, e.g., for model merging, multi-objective optimization, and privacy-preserving machine learning25, 26. Here, we explore the untapped potential of task arithmetic as a new transfer learning strategy for molecule design. We first show how molecular task arithmetic unlocks zero-shot molecule design, i.e., generating positive molecules without any positive labeled molecules. Next, we use molecular task arithmetic for dual-objective tasks, and as a drop-in replacement for the traditional supervised finetuning in few-shot applications. Across 20 studied design tasks, molecular task arithmetic surpasses supervised finetuning by creating more diverse and accurate molecules. Molecular task arithmetic is simple, applicable to all models optimized via gradient descent, and bears the potential to be the standard transfer learning pipeline for molecule design. Our key contributions are as follows:
  • We propose a simple and widely-applicable transfer learning strategy – molecular task arithmetic. Molecular task arithmetic tackles the inherent data limitations in de novo design by harnessing negative molecules, a direction explored for the first time here. Across experiments, models trained with molecular task arithmetic design more accurate and diverse molecules.
  • We unlock zero-shot ligand-based de novo design with transfer learning. Since molecular task arithmetic uses only negative molecules to learn task directions, it can be used even when no positively labeled data is available. Models trained with molecular task arithmetic can design molecules fulfilling one or multiple target properties, with minimal side effects on non-targeted properties.
@@
  • We improve few-shot molecule design with traditional transfer learning pipelines. When used to augment the traditional finetuning based on positive molecules, molecular task arithmetic increases the number of diverse and desirable designs across 20 experiments. This shows the potential of molecular task arithmetic to replace the dominant transfer learning practices, by allowng to leverage both positive and negative molecules simultaneously.
Fig. 1: Molecular task arithmetic. (a) SMILES string creation from a molecular graph. Atoms are annotated by the periodic table symbols and double bonds are denoted with ‘=’ sign. Lowercase symbols describe aromatic bonds and brackets annotate branches. (b) Chemical language modeling. A next token prediction task can be defined on SMILES strings using a sequence model such as an LSTM. The learned language model can be sampled for molecule design. (c) Zero-shot design. Molecular task arithmetic learns a task direction in the model weight space by applying finetuning on negative molecules (dashed purple arrow). The learned direction is traversed in the opposite direction (solid orange arrow) for zero-shot molecule generation. (d) Dual-objective design. By learning multiple task directions independently, molecular task arithmetic can train multi-objective models. (e) Few-shot design. Finetuning the zero-shot model with known positive molecules (dashed orange arrow) enables molecular task arithmetic to use all available data for the task.
2 Background
Task arithmetic
@@ 2 Background
In the weight space of a trained model, certain directions correspond to improvements in performance on specific tasks 25. These directions are known as “task vectors”25, 27, 28, 29. In a transfer learning setting, where data for a task $t$ are available, a task vector $\tau^{t}$ can be computed by subtracting the pretrained model weights from the finetuned model weights25:
@@
where $\lambda$ is a scaling factor that determines the step size in the direction. Termed as “forgetting via negation”25, this approach can remove undesired behavior from deep models, such as toxic language generation25. Task vectors can also be added together to create models with multi-task capabilities (“learning via addition”25):
@@
where $\tau^{t_{i}}$ is the task vector learned for the task $t^{i}$. Both in Eq. (2) and in Eq. (3), $\lambda$ controls how far the final model lands from the pretrained one in the weight space. While too large $\lambda$ values can have detrimental side effects on model performance, too low values fail to sufficiently edit the model behavior.
Molecular string generation for de novo design
Molecular strings are among the most popular molecular representations for de novo molecule design 30, 31, 32, 33. The most well-established representation is the Simplified Molecular Input Line Entry System (SMILES 31) strings (Fig. 1a). SMILES strings are created by traversing a molecular graph in any chosen direction, starting from any non-hydrogen atom, and annotating atomic symbols, plus branches and bond types. Multiple SMILES strings (often referred to as randomized SMILES) can be created by starting from different atoms (and/or with different traversals); this characteristic allows for data augmentation, where multiple instances referring to the same molecule are used to inflate the amount of training data34. Canonicalization algorithms have been proposed to obtain a single SMILES string from any molecule (i.e., by standardizing the starting atom and the graph traversal rules 35, 36). Canonicalization can aid in duplicate molecule detection and molecule ranking 37, 38, 39. Both versions are commonly used for molecule design, with different advantages and drawbacks 40.
For modeling purposes, the SMILES string represents a molecule $m$ as $\{s_{1},s_{2},\dots,s_{n}\}$, where $s_{i}$ corresponds to the $i^{th}$ token in the molecular string. The next token prediction language modeling objective is then defined as: $\max\sum_{m}^{M}\sum_{i}\log P(s_{i}|s_{<i};\theta)$, where $\theta$ is typically a long-short term memory network41, 42, 14 or a transformer43, 44, 15 (Fig. 1b), and $M$ is a dataset of SMILES strings.
In a transfer learning setting, two phases typically occur: (a) pretraining, where $M$ is selected to contain hundreds of thousands of molecules to learn key elements, such as generating ‘chemically valid’ SMILES, and to learn basic molecular properties17; and (b) (self-)supervised finetuning, where a curated set of molecules with desired properties is used, with the same training objective to condition the model45, 46, 17. New molecules can then be designed by sampling the learned multinomial distribution autoregressively. Chemical language models have already demonstrated powerful design capabilities for drug discovery and are a well-established paradigm for molecule design 47, 48, 4, 49, 40, 14, 15, 16.
@@ 3 Molecular Task Arithmetic
Why task arithmetic? Chemical space is vast but sparse: it is estimated to contain 1060 molecules, but molecules of desirable properties are rare 3. Current training strategies rely on these rare, ‘positive’ molecules (i.e., molecules possessing one or more desirable properties), and therefore, the limited data availability and diversity form a bottleneck. Here we ‘look the other way’ and propose a new transfer learning strategy that can leverage the abundant and diverse ‘negative’ molecules, that is, molecular task arithmetic.
Molecular task arithmetic views “forgetting via subtraction” (Eq. 2) as an opportunity to learn only from negative data to design positive molecules. For the task $t$ of designing molecules with a desirable property, molecular task arithmetic finetunes a pretrained model ($\theta_{pt}$) with molecules having non-desirable values of this property. This allows obtaining $\theta_{opp}$, a model that can generate negative molecules. Then the task vector $\tau^{t}$ is computed by subtracting $\theta_{pt}$ from $\theta_{opp}$. Finally, a model that creates the molecules of positive property, $\theta_{task}$ (Fig. 1c), can be constructed by subtracting $\tau^{t}$ from $\theta_{pt}$ with a scaling factor $\lambda$:
@@
In other words, molecular task arithmetic learns the ‘direction’ of a molecular property using negative molecules and then ‘looks the other way’ – by moving the pretrained model in the opposite direction to design positive molecules. By not using any labeled positive molecules, this formulation opens the doors for unexplored tasks, e.g., zero-shot molecule design.
To test the impact of task arithmetic on steering de novo design of bespoke molecules, we have applied it to easily computable properties that are relevant for drug discovery, namely: (a) fraction of sp3-hybridized carbons (frac. sp3C), which reflects molecular complexity and 3D characteristics; (b) octanol-water partitioning coefficient (logP), influencing membrane permeability and solubility; (c) number of hydrogen bond donors, which affects a molecule’s ability to form key interactions with biological targets; (d) number of rings (No. Rings), linked to molecular rigidity and target binding specificity; and (e) topological surface area (TPSA).
We test molecular task arithmetic on chemical language models with an LSTM backbone. We pretrained the LSTM on 1.5M drug-like molecules and define conditional molecule design experiments as designing molecules with a higher or lower value than a predefined threshold of the five selected molecular properties (Fig. A1). By defining tasks for randomized and canonical SMILES and per property, we create 20 design tasks. Five splits of 1024 train, 256 validation, 256 test molecules, and 10240 negative molecules are curated per task for finetuning and molecular task arithmetic (see Appendix for more details).
4 Experiments
@@ 4.1 Can Task Arithmetic Even Generate Valid Molecules?
The language of SMILES strings possesses a strict grammar to represent molecules, where single-character changes might yield ‘chemically invalid’ strings (i.e., strings that cannot be converted back into molecules). Hence, our first question was whether the capacity to generate chemically valid strings could be retained under the model weight manipulation introduced by task arithmetic.
To answer this question, we applied molecular task arithmetic with $\lambda=0.50$ (Eq. 3) across the ten single-objective design tasks and the two SMILES formats (canonical and randomized). We then computed molecular validity (defined as the ratio of unique and chemically valid design strings among 100,000 generations), as a well-established control of the capacity of a model to learn the ‘chemical language’ of SMILES. Molecular task arithmetic obtains an average validity of 79.0% with randomized SMILES strings and 59.2% with canonical SMILES strings. While these values drop from the pretrained models (which were optimized for validity, obtaining values of 95.2% and 95.8% for randomized and canonical SMILES strings, respectively), molecular task arithmetic preserves its ability to generate chemically meaningful strings, especially with randomized SMILES strings. Remarkably, this finding indicates that the weight space of a chemical language model can be traversed with no loss guidance – opening new and unexplored avenues for de novo molecule design with generative deep learning.
In what follows, we explore the benefits and limitations of traversing the model weight space with molecular task arithmetic, with the following objectives:
  1. 1.
    Exploring zero-shot de novo design, where only molecules with ‘undesired’ properties are used to design ‘desirable’ molecules. In this context, we explore single-property molecule design (Fig. 1c), and dual-objective molecule design (Fig. 1d), where single and pairs of properties are targeted, respectively.
  2. 2.
    Understanding the effect of task vector scaling, where we analyze how the magnitude of the molecular task vector ($\lambda$) affects the capacity to explore the chemical space in search for molecules with desirable properties.
  3. 3.
    Boosting few-shot molecule design (Fig. 1e), by augmenting traditional finetuning (which relies on positive molecules only) with negative molecules, via molecular task arithmetic.
@@ 4.2 Zero-Shot Molecule Design
Fig. 2: Zero-shot molecule design with molecular task arithmetic. Models were trained on 10 design tasks of randomized SMILES strings, with molecular task arithmetic and supervised finetuning. 100,000 molecules were designed and clustered. (a) The number of cluster centers that possess the desired property, (b) the ratio of the designs that satisfy the design task, and (c) number of clusters were computed. The pretrained model is also included in the analysis as a baseline. Bar heights report the mean statistics across five training sets, and error bars denote the standard deviation. Finetuning was conducted up to the training sets of 1024 molecules, while molecular task arithmetic used 10,240 negative molecules and no labeled positive data.
@@
We benchmarked molecular task arithmetic with (a) supervised finetuning and (b) pretraining only. We experimented with varying values of $\lambda$ (controlling the deviation from the pretrained model, Eq. 3), in particular, between 0.10 and 1.00 with a step size of 0.05. We experimented with fractions of the finetuning dataset (composed of 1024 molecules) ranging from 1% to 100%, with increments of 10%. For task arithmetic, 10,240 negative molecules were used. We evaluated the models by the number of successful design clusters, i.e., the number of design clusters that possess the desired property. This metric combines the evaluation of diversity and accuracy; the higher, the better. The highest number of successful design clusters obtained across $\lambda$ values (task arithmetic), or training subsets (supervised finetuning) was reported for each property. In general, models trained on randomized SMILES yielded more successful design clusters than their canonical counterparts (Fig. A2). Due to the better performance of randomized SMILES, in what follows, we report only the results for randomized SMILES for the zero-shot experiments, while results for canonical SMILES strings were included in the Appendix.
@@
Across design tasks, molecular task arithmetic yielded a higher number of successful design clusters than pretraining and finetuning (9.5K and 11.7K more on average, respectively; Fig. 2a and Fig. A3a ). Except for the $\uparrow$ No. Rings’ task with randomized SMILES, the performance gain by molecular task arithmetic over supervised finetuning is statistically significant across datasets (Mann-Whitney U test, p-value $<$ 0.01). This is a striking finding: despite using no positively labeled molecules, molecular task arithmetic can design a higher number of successful diverse molecules than models fine-tuned on up to 1024 positive compounds. This demonstrates the potential of task arithmetic for generative drug discovery without any positively labeled molecular examples.
@@
To further shed light on the success of molecular task arithmetic, we inspect the success rate (ratio of designs satisfying the task; Fig. 2b and Fig. A3b) and number of total clusters (Fig. 2c and Fig. A3c). Results show that molecular task arithmetic can create, on average, 24.4K more diverse molecule designs than the fine-tuned models and 7.4% more accurate designs than pretraining, while finetuning outperforms molecular task arithmetic by 14.1% in success rate. This is, on one hand, expected, since finetuning has explicit access to molecules fulfilling the desired property. On the other hand, it is surprising that molecular task arithmetic can still create more desired cluster centers with less successful designs. Taken together, these findings display that molecular task arithmetic can preserve the diversity gained by pretraining during conditioning, while finetuning ‘forgets’. Forgetting and loss of generalization with finetuning are observed in different deep learning contexts50, 51, 52, 53, 54, 55, and here we show that task arithmetic can be a solution for molecule design, to maintain higher diversity while achieving accuracy.
Fig. 3: Side effects of transfer learning strategies. Kolmogorov-Smirnov distance between the molecular properties of the designs on SMILES tasks and the pretraining set is computed for off-target properties. Lower distance to the pretraining set indicates that the conditioning caused less change in non-conditioned molecular properties. Bar heights display the means across five training splits, and error bars display the standard deviations.
Side effects.
Another desirable trait of model conditioning is the possibility of having minimal side effects on the ‘off-target’ properties. Being able to minimize side effects would allow for a more precise steering of the model in the chemical space, by preserving desirable molecular properties (i.e., synthesizability, lack of toxicity, etc.). We quantify the side effects of both finetuning and molecular task arithmetic for all design tasks, on the remaining four descriptors. These descriptors have a low correlation with the ‘target’ properties (Table A3). For each task, we computed the Kolmogorov-Smirnov (KS) distance between the off-target molecular properties of the designs and the pretraining set (the lower the KS, the more similar the distributions). The usefulness of this analysis is corroborated by observing the KS distances of the designs obtained by pretraining (Fig. A4), which are all below 5%.
Across the ten design tasks for randomized SMILES, molecular task arithmetic showed a smaller mean distributional distance to the pretraining set in 32 of the 40 comparisons (Fig. 3), of which 15 were confirmed by a statistical test (Mann-Whitney U test, p-value $<$ 0.01). On canonical SMILES strings, all models have a higher average distance (Fig. A5), possibly because canonicalization restricts the models to explore ‘side-effect-free’ portions of the chemical space. Overall, the analysis shows that molecular task arithmetic can combine its accuracy and diversity on the design task, with less side-effect than the traditional finetuning approach. This suggests that task arithmetic allows for a more controlled steering of the models in the desirable portions of the chemical space.

### v2

@@ Abstract
The scarcity of molecules with desirable properties (i.e., ‘positive’ molecules) is an inherent bottleneck for generative molecule design. To sidestep such obstacle, here we propose molecular task arithmetic: training a model on diverse and abundant negative examples to learn ‘property directions’ – without accessing any positively labeled data – and moving models in the opposite property directions to generate positive molecules. When analyzed on 33 design experiments with distinct molecular entities (small molecules, proteins), model architectures, and scales, molecular task arithmetic generated more diverse and successful designs than models trained on positive molecules in general. Moreover, we employed molecular task arithmetic in dual-objective and few-shot design tasks. We find that molecular task arithmetic can consistently increase the diversity of designs while maintaining desirable complex design properties, such as good docking scores to a protein. With its simplicity, data efficiency, and performance, molecular task arithmetic bears the potential to become the de-facto transfer learning strategy for de novo molecule design.
@@ 1 Introduction
Discovering one drug molecule can take over a decade and billions of dollars wouters2020estimated. The first obstacle is charting the ‘chemical space’ effectively bohacek1996art. Chemical space is estimated to contain approximately 1060 drug-like molecules, with a scarcity of molecules possessing desirable properties, e.g., bioactivity towards a pharmaceutically relevant target. Generative deep learning has emerged as a revolutionary technology for drug discovery – with the potential to shorten de novo design pipelines from years to weeks wu2024tamgen, ghazi2025quantum.
Generative drug discovery faces an inherent challenge: molecules with desirable properties (e.g., bioactivity) are not only scarce but also may lack structural diversity. In the case of early-stage drug discovery, positive molecules (or ‘hits’) can be as rare as 1% in large, diverse molecule libraries. For new pharmacological targets, identifying bioactive molecules is often time-consuming, with the vast majority of candidates proving inactive compared to the few that show activity. Finally, bioactive molecules are often optimized through minor structural edits to explore structure-activity relationships, resulting in medicinal chemistry datasets typically containing a few hundred molecules with low structural diversity sun2017excape, tran2020lit, van2022exposing. Transfer learning has served to mitigate data scarcity by (a) pretraining on large unlabeled molecular sets, then (b) finetuning on molecules with desirable properties (e.g., bioactivity) segler2018generating. Despite this, the constraints of drug discovery data – where active compounds are vastly outnumbered by inactive ones, or might not be available – can still limit model effectiveness. In this context, leveraging the large wealth of ‘negative’ data is an unexplored avenue to boost the potential of generative deep learning for drug discovery. Stemming from this observation, this work introduces task arithmetic for the first time into the molecular sciences, as an effective strategy for chemical space exploration leveraging negative data.
Task arithmetic manipulates model weights to transfer or combine learned properties across tasks ilharco2023editing. It has shown promise in vision and language applications, such as model merging and multi-objective optimization ilharco2023editing, choi2024revisiting. Here, we apply it to an intriguing challenge: training with negative data to steer models away from undesirable chemical regions in order to generate positive molecules. This is an especially daunting task because (a) chemical space is vast, discrete, and sparse, and (b) structure-property relationships are non-linear. Here, we investigate whether task arithmetic can open new opportunities to efficiently chart the dark chemical universe. Our large-scale and systematic study across 33 (bio)chemical design tasks shows that it can design positive molecules by using negative ones, thereby remarkably advancing the capabilities of generative deep learning to navigate the chemical space.
The key contributions of this work are the following:
  • We introduce molecular task arithmetic (MTA) in generative drug design for the first time, as a strategy to overcome data limitations typical of drug discovery and leverage negative molecules in a novel way.
  • We show that MTA enables zero-shot ligand-based de novo design, by designing molecules fulfilling one or multiple target properties starting from negative molecules, even in cases where no positive molecules were available to the models.
@@
  • We augment few-shot molecule design with task arithmetic. When used to augment traditional transfer learning pipelines, MTA increases the number of diverse and desirable designs across tasks of increasing difficulty.
Our results display the potential of molecular task arithmetic to replace dominant ligand-based de novo design practices, by leveraging molecules with undesirable properties in an unprecedented way.
Fig. 1: Molecular task arithmetic (MTA). (a) Zero-shot design. Molecular task arithmetic learns a task direction in the model weight space by finetuning on negative molecules (dashed purple arrow). The task vector is traversed in the opposite direction (solid orange arrow). (b) Multi-objective design. Multiple task directions are learned independently, and then combined. (c) Few-shot design. Task arithmetic is applied on the model finetuned with negatives, and then known positive molecules are used (dashed orange arrow).
2 Background and Related Work
Task arithmetic.
@@ 2 Background and Related Work
Certain directions in the weight space of a trained model – task vectors – correspond to a change in performance on specific tasks ilharco2023editing, yang2025task. In a transfer learning setting, where data for a task $t$ are available, a task vector $\tau^{t}$ can be computed by subtracting the pretrained model weights from the finetuned model weights ilharco2023editing:
@@
where $\lambda$ is a scaling factor that determines the step size in the direction. Termed as “forgetting via negation”, this approach can remove undesired behavior from deep models, such as toxic language generation ilharco2023editing. Task vectors can also be added together to create models with multi-task capabilities – “learning via addition” ilharco2023editing:
@@
where $\tau^{t_{i}}$ is the task vector learned for the task $t^{i}$. In Eq. 2 and 3, $\lambda$ controls how far the final model lands from the pretrained one in the weight space. While too large $\lambda$ values have detrimental side effects on model performance, too low values fail to edit the model behavior sufficiently.
Molecule design with deep learning.
Early deep learning for molecule design approaches trained sequence models on string representations of molecules with a next-token prediction language modeling task olivecrona2017molecular, segler2018generating. Recent works used molecular graphs and point clouds to represent molecules, combined with graph neural networks and diffusion xia2019graph, wang2025diffusion. While these approaches enable more fine-grained representation and higher steerability in learning, decoder-only language models of molecular strings, termed chemical language models, remain popular to date due to their simplicity and performance grisoni2023chemical. SMILES (Fig. A1a) is the most popular string representation for chemical language modeling. In a transfer learning setting, two phases typically occur: (a) pretraining, a sequence model is trained with a language modeling task (Fig. A1b), on several millions of molecules to learn key elements, e.g., generating ‘chemically valid’ SMILES and basic molecular properties; and (b) (self-)supervised finetuning, where a curated set of molecules with desired properties is used, with the same training objective to condition the model skinnider-2021. New molecules can then be designed by sampling the learned multinomial distribution autoregressively. This pipeline has become a well-established standard merk2018novo, wu2024tamgen, ghazi2025quantum.
@@ 3 Molecular Task Arithmetic
Why task arithmetic? Chemical space is vast but sparse: it is estimated to contain 1060 molecules, but molecules of desirable properties are rare. Current finetuning strategies rely on these rare, ‘positive’ molecules and therefore, the limited data availability and diversity form a bottleneck. Here we ‘look the other way’ and propose a new transfer learning strategy that aims to leverage the abundant and diverse ‘negative’ molecules, that is, molecular task arithmetic (MTA).
MTA views “forgetting via subtraction” (Eq. 2) as an opportunity to learn only from negative data to design positive molecules. For the task $t$ of designing molecules with a desirable property, MTA finetunes a pretrained model ($\theta_{pt}$) with molecules having non-desirable values of this property. This yields a model ($\theta_{opp}$) that can generate negative molecules. The task vector $\tau^{t}$ is then computed by subtracting $\theta_{pt}$ from $\theta_{opp}$. Finally, a new model ($\theta_{task}$) can be constructed by subtracting $\tau^{t}$ from $\theta_{pt}$ with a scaling factor $\lambda$ (Fig. 1a):
@@
Our main scientific question is whether $\theta_{task}$ can generate chemically-valid molecules with desirable properties. In other words, whether molecular task arithmetic can learn the ‘direction’ of a molecular property using negative molecules and then ‘look the other way’ – by moving the pretrained model in the opposite direction to design positive molecules. By not using any labeled positive molecules, if successful, task arithmetic might open doors to unexplored tasks, e.g., zero-shot ligand design.
4 Experimental Settings
Physico-chemical properties.
As a first systematic analysis, we applied MTA to five physico-chemical properties that are relevant for drug-likeness, and at the same time are easy and fast to compute: (a) fraction of sp3-hybridized carbons (frac. sp3C); (b) octanol-water partitioning coefficient (logP); (c) number of hydrogen bond donors; (d) number of rings (No. Rings); and (e) topological surface area (TPSA). We defined conditional molecule design experiments as designing molecules with a higher or lower value than a predefined threshold of the five selected molecular properties (Fig. A2), for a total of 10 design tasks. For each task, we curated five training, validation and test splits (containing 1024, 256 and 256 positive molecules, respectively). Additionally, we defined three dual-objective tasks: (a) high fraction of sp3-hybridized carbons and a high number of hydrogen bond donors, (b) high lipophilicity and low TPSA, and (c) high number of rings and low TPSA. We selected minimally correlated properties (absolute correlation lower than 0.3; Table A1) to ensure neither conflicting nor trivial tasks.
Ligand-target docking.
We applied MTA to achieve a more challenging task: designing molecules that dock well into a protein target using negative data. We chose three pharmacologically relevant proteins: (a) coagulation factor II (F2), (b) glucocorticoid receptor (NR3C1), and (c) peroxisome proliferator-activated receptor delta (PPARD). We curated (a) binding molecules with good docking scores for supervised finetuning and (b) non-binding molecules with poor docking scores for molecular task arithmetic (see Appendix for details). Both well-docking and poorly-docking molecules constitute less than 5% of the pretraining set, further increasing the difficulty of the task (Fig. A3).
Overall setup.
We relied on chemical language models (long-short term memory networks, LTSMs, 3.17M parameters) pretrained on 1.5M SMILES strings from ChEMBL gaulton2017chembl. For any task, we (a) finetuned models with positive molecules, and (b) trained molecular task arithmetic models with negative molecules. What follows is structured based on the following objectives:
  1. 1.
    Generating ‘chemically valid’ molecules, where we evaluate if the model weight manipulation introduced by task arithmetic destructs generative capabilities of the model.
  2. 2.
    Achieving zero-shot de novo design, where only molecules with ‘undesired’ properties are used to design ‘desirable’ molecules, both for single- and dual-objective design (Fig. 1a,b).
  3. 3.
    Understanding the effect of task vector magnitude, where we analyze how scaling the molecular task vector via $\lambda$ affects chemical space exploration for desirable molecules.
  4. 4.
    Boosting few-shot molecule design (Fig. 1c), by augmenting traditional finetuning (which relies on positive molecules only) with negative molecules, via molecular task arithmetic.
  5. 5.
    Designing well-docking molecules, where we study if molecular task arithmetic can harness poor-docking molecules to design well-docking ones, as a proxy for designing bioactive molecules starting from inactive ones.
Moreover, as an additional modality, we applied molecular task arithmetic to designing highly-ordered proteins. This serves to study if molecular task arithmetic can be extended to additional molecular entities and model architecture (large-scale pretrained transformer, ferruz2022protgpt2) to design proteins with fewer disordered regions, starting from disordered proteins.
5 Results
@@ 5.1 Can Task Arithmetic Even Generate Valid Molecules?
The language of SMILES strings possesses a strict grammar to represent molecules, where single-character changes might yield strings that cannot be converted back into molecules, i.e., invalid SMILES strings. Hence, our first question was whether the capacity to generate valid SMILES is retained under the model weight manipulation introduced by task arithmetic. To answer this question, we applied molecular task arithmetic (MTA) with $\lambda=0.50$ (Eq. 3) across the 20 single-objective design tasks, consisting of two SMILES formats. We then computed validity (the ratio of valid and unique SMILES strings among 100,000 generations), as a well-established measure of a model’s capacity to learn SMILES syntax. Molecular task arithmetic obtains an average validity of 79.0% with randomized SMILES strings and 59.2% with canonical SMILES strings. While these values drop from the pretrained models (which were optimized for validity, obtaining values of 95.2% and 95.8% for randomized and canonical SMILES strings, respectively), MTA preserves its ability to generate valid molecules, especially with randomized SMILES strings. Remarkably, this finding indicates that the weight space of a chemical language model can be traversed with no loss guidance – opening unexplored avenues in generative deep learning for molecule design.
@@ 5.2 Zero-Shot Molecule Design
We benchmarked MTA with (a) supervised finetuning and (b) pretraining only. We experimented with varying values of $\lambda$ (controlling the deviation from the pretrained model, Eq. 3), in particular, between 0.10 and 1.00 with a step size of 0.05. We experimented with fractions of the finetuning dataset (composed of 1024 molecules) ranging from 1% to 100%, with increments of 10%. For task arithmetic, 1024 negative molecules (finetuning set of the opposite property) were used. We evaluated the models by the number of successful design clusters, i.e., number of structurally-distinct molecular clusters identified by rdkit’s LeadPicker module with a preset distance threshold among the designs with the desired property. This metric combines the evaluation of diversity and accuracy; the higher, the better. The highest number of successful design clusters obtained across $\lambda$ values (task arithmetic), or training subsets (supervised finetuning) was reported for each property. In general, models trained on randomized SMILES yielded more successful design clusters than their canonical counterparts (Fig. A4) as also observed elsewhere arus2019randomized. Due to the better performance of randomized SMILES, in what follows, we report only the results for randomized SMILES and include results for canonical SMILES strings in the Appendix.
Fig. 2: Zero-shot single-objective molecule design. Molecular task arithmetic vs fine-tuning, across 10 design tasks (100,000 molecular designs). The pretrained model is included as a baseline, and average and standard deviation are reported. (a) number of cluster centers that possess the desired property; (b) ratio of designs that satisfy the design task; (c) number of clusters.
@@
Across design tasks, MTA yielded a higher number of successful design clusters than pretraining and finetuning (9.5K and 11.7K more on average, respectively; Fig. 2a and Fig. A5a ). Except for the $\uparrow$ No. Rings’ task with randomized SMILES, the performance gain by MTA over supervised finetuning is statistically significant across datasets (Mann-Whitney U test, p-value << 0.01). This is a striking finding: despite using no positively labeled molecules, molecular task arithmetic can design a higher number of successful diverse molecules than models fine-tuned on up to 1024 positive compounds. This demonstrates the potential of task arithmetic for generative drug discovery without any positively labeled molecular examples.
@@
To further shed light on the success of MTA, we inspect the success rate (ratio of designs satisfying the task; Fig. 2b and Fig. A5b) and number of total clusters (Fig. 2c and Fig. A5c). Results show that MTA creates, on average, 24.4K more diverse molecule designs than the finetuned models and 7.4% more accurate designs than pretraining, while finetuning outperforms MTA by 14.1% in success rate. This is, on one hand, expected, since finetuning has explicit access to molecules fulfilling the desired property. On the other hand, it is surprising that MTA can create more clusters with less successful designs. Together, these findings display that MTA can preserve the diversity gained by pretraining during conditioning, while finetuning ‘forgets’. Forgetting with finetuning is observed in different deep learning contexts vieira2024much, lampinen2025generalization, and here we show that MTA is a technique that can maintain higher diversity while conditioning molecule designs.
Distribution shifts.
To study how each training strategy shifts the property distribution, we first assess extrapolation capabilities for target properties, i.e., whether they can design molecules with property values outside the training sets. We inspected high-value design tasks with randomized SMILES and computed the maximum value of the task properties for the pretraining set and the designs (Table A2). MTA designs possessed values equal to the theoretical maximum or above the pretraining values in four of five cases (except for logP). In contrast, finetuning could saturate only one design task (Frac. sp3C). This finding shows that molecular task arithmetic can extrapolate beyond the property distributions on which it was trained, unlike supervised finetuning, offering potential to explore new portions of the chemical space.
We next study how MTA and finetuning impact ‘off-target’ properties. We computed the distributional distance between off-target properties of the designs and the pretraining set. MTA designs possessed a smaller distance to the pretraining set than finetuning ones in general (Fig. A6, A7, A8), showing that MTA enables a more controlled steering of the models in the desirable portions of the chemical space (more details in the Appendix).
Out-of-distribution generalization.
How does molecular task arithmetic perform when no positive molecule is available in the pretraining set? Here we seek an answer to this question via six design tasks across three descriptors. We picked fraction of sp3C, logP, and TPSA, and pretrained three LSTMs (0.41M parameters) on 250K molecules from ChEMBL, such that (a) 0.2 $\leq$ Frac. sp3C $\leq$ 0.85, (b) 1 $\leq$ logP $\leq$ 6, and (c) 50 $\leq$ TPSA $\leq$ 100, respectively. We defined design tasks as designing molecules whose properties are outside the pretraining set limits. We applied MTA on the molecules of the ‘negative’ task ($0.10\leq\lambda\leq 1.00$; step size 0.10), i.e., training sets containing no molecules possessing the desired property. Fine-tuning with ‘positive’ molecules was also performed to assess performance when positive data are available. The number of successful design clusters was measured (with 100K designs) in increasing dataset sizes from 10 to 1024, across five repeats with different negative/positive molecule sets.
h Refer to caption
Fig. 3: Out-of-distribution design. The number of successful design clusters was computed at increasing sizes of ‘negative’ and ‘positive’ training sets. Mean (solid lines) and standard deviation (shaded areas) are reported (100,000 designs, five training-validation splits).
MTA designs 10K-50K successful molecule clusters across six design tasks (Fig. 3), although it is neither pretrained nor finetuned on any positive molecules. This represents a remarkable increase compared to the 1K-5K clusters obtained by the pretrained model, demonstrating the performance boost of MTA even in OOD settings. When 1024 positive molecules are available, finetuning achieves up to 55K successful design clusters. This highlights the added value of positive data in pushing the pretrained model toward a new region of the property distribution; MTA on negative data constitutes a valuable alternative. In fact, our analysis shows approximate “positive-to-negative data exchange rates“, which depend on the target property. For ‘Frac. sp3C’ and ‘logP’, even ten positive molecules are enough to outperform 1024 negative molecules, while for $\uparrow$ TPSA’, at least 250 positive molecules are needed to match the performance of ten negative ones. For $\downarrow$ TPSA’, 50 negative molecules can yield the same number of successful design clusters as 500 positive molecules, demonstrating that the value of negative data can even be higher than positive molecules in certain low-data settings. Taken together, our results show that MTA can design positive molecules via negative ones, even when no positive molecules exist in the pretraining set. Positive data remains, in general, valuable for out-of-distribution generalization (whenever available), while negative data can yield better performance for certain tasks in the low-data scenarios.